A woman can read the large print in a newspaper only when it is at a distance of 65 cm or more from her eyes.


What should be the refractive power (in diopters) of her glasses (worn 1.9 cm from the eyes), so she can read the newspaper at a distance of 25 cm from the eyes?
Answer #1

Solution)

Here,

u = 25 cm -1.9 cm = 23.9 cm

Also,

v = -65+1.9=- 63.1 cm

Applying len's equation, we have

1/v+1/u =1/f

1/f =1/-63.1 cm + 1/22.9

f=35.94 cm

Hence,

Power = 100/f = 100/35.94=+2.78 D

========